OpenStudy (anonymous):

what is the solution of the system of equations? 2x+2y+3z=-1 4x+4y+3z=-5 4x+2y+2z=0

OpenStudy (dumbcow):

you want the long way or the short cut using matrices and a calculator

OpenStudy (anonymous):

showing the work way, guess its long way

OpenStudy (dumbcow):

haha ok i recommend using a bunch of elimination to isolate a single variable lets start by eliminating x: pick 2 equations and eliminate x leaving an equation with y and z 4x +4y +3z = -5 -4x -2y -2z = 0 (I multiplied by -1 so x's would cancel) --------------- 2y +z = -5 Repeat with 2 other equations -4x -4y-6z = 2 (I multiplied by -2 so x's would cancel) 4x +4y +3z =-5 --------------- -3z = -3 (just so happens y cancels as well, saves a step) z = 1 substitute it into previous equation 2y + 1 = -5 y = -3 substitute both z,y values into one of original equation to get x 2x +2(-3) +3(1) = -1 2x -6 +3 = -1 x = 1 Solution: x=1,y=-3, z=1